two suitcases are similar rectangular solids. The length of the second suitcase is twice the length of the first. The first suitcase has a volume of 1568 cubic inches what is the volume of the second suitcase in cubic inches.
Is there a formula for this. My son and I have tried and tried and just can't do this. Thank you

Answer by ankor@dixie-net.com(22740)

Two suitcases are similar rectangular solids.
The length of the second suitcase is twice the length of the first.
The first suitcase has a volume of 1568 cubic inches.
What is the volume of the second suitcase in cubic inches?
:
Since the suitcases are similar, we can assume the width and height of the second suitcase, is also twice the first;
:
First suitcase
L * W * H= 1568
:
2nd suitcase
2L * 2W * 2H = 8*L*W*H
:
Replace L*W*H with 1568
8(1568) = 12544 cu/in